Xavier has been working at his first post-college job for almost a year when his company gives him a raise, resulting in a paycheck increase of $200, for a total of $400 extra in take-home pay every month. He makes a quick list of possible ways to use that money, along with relevant notes for each.


1
Expert's answer
2021-02-11T12:44:26-0500

Task#162720 solution

Option 1

Xavier can maintain his monthly budget as it is and save the extra $200 at the end of every month for like 5 years at 8 % interest rate per annum and use the money wisely based on his future dreams.

After 5 years, he will have:

Future value (FV) = Monthly saving (PMT)*Future value interest factor (FVIFA), period 5 years, 7% rate where;

"FV=PMT\times FVIFA"


"FVIFA= \frac{(1+ r) ^n-1}{r}, n= 5\times12=60, r= \frac{0.07}{12}"

"FV= 200\times PVIFA"


"FV=200\times{(1+\frac{0.07}{12})^60-1 \over \frac{0.07}{12}}"


"FV= 200\times71.5884782=14,317.70"

Future value= $14,317.70


He can use this amount to buy a house or any other asset after five years.


Option 2. Increase his current budget/spending

Xavier can also increase his current budget and live a more comfortable life using the salary increment.


Option 3. Purchase an asset which will give him a chance to pay a monthly premium within the increment.

Xavier can opt to buy an asset and pay monthly using the salary increment


Option 4. Do nothing

He may as well do nothing and leave the money in his account for precautionary, speculative or transactionary motives